# A Generalized Analytical Framework for the Nonlinear Best-Worst Method

### v1

@@ 4.1.1 Calculation of Weights in the Presence of Unique Best and Worst Criterion
Now, assume that $\eta^{*}=\epsilon_{i_{0},j_{0}}\text{ for some }i_{0},j_{0}\in D$. As discussed in Proposition 1, we get $0\leq\eta_{bi_{0}},\eta_{i_{0}w},\eta_{bj_{0}},\eta_{j_{0}w}\leq\eta^{*}=\epsilon_{i_{0},j_{0}}$ such that $\tilde{a}_{bi_{0}}^{*}\in\{a_{bi_{0}}-\eta_{bi_{0}},a_{bi_{0}}+\eta_{bi_{0}}\}$, $\tilde{a}_{i_{0}w}^{*}\in\{a_{i_{0}w}-\eta_{i_{0}w},a_{i_{0}w}+\eta_{i_{0}w}\}$, $\tilde{a}_{bj_{0}}^{*}\in\{a_{bj_{0}}-\eta_{bj_{0}},a_{bj_{0}}+\eta_{bj_{0}}\}$ and $\tilde{a}_{j_{0}w}^{*}\in\{a_{j_{0}w}-\eta_{j_{0}w},a_{j_{0}w}+\eta_{j_{0}w}\}$. Since $(\tilde{A}_{b}^{*},\tilde{A}_{w}^{*})$ is consistent, we have $\tilde{a}_{bi_{0}}^{*}\times\tilde{a}_{i_{0}w}^{*}=\tilde{a}_{bj_{0}}^{*}\times\tilde{a}_{j_{0}w}^{*}$. Now, it is easy to observe that if any one of $\eta_{bi_{0}}$, $\eta_{i_{0}w}$, $\eta_{bj_{0}}$ or $\eta_{j_{0}w}$ is strictly less than $\epsilon_{i_{0},j_{0}}$, then at least one of the remaining three must be strictly greater than $\epsilon_{i_{0},j_{0}}$, which is not possible. This gives $\eta_{bi_{0}}=\eta_{i_{0}w}=\eta_{bj_{0}}=\eta_{j_{0}w}=\epsilon_{i_{0},j_{0}}$. It is clear that the only combination among the sixteen possible values of $(\tilde{a}_{bi_{0}}^{*},\tilde{a}_{i_{0}w}^{*},\tilde{a}_{bj_{0}}^{*},\tilde{a}_{j_{0}w}^{*})$ that satisfies $\tilde{a}_{bi_{0}}^{*}\times\tilde{a}_{i_{0}w}^{*}=\tilde{a}_{bj_{0}}^{*}\times\tilde{a}_{j_{0}w}^{*}$ is $(a_{bi_{0}}+\epsilon_{i_{0},j_{0}},a_{i_{0}w}+\epsilon_{i_{0},j_{0}},a_{bj_{0}}-\epsilon_{i_{0},j_{0}},a_{j_{0}w}-\epsilon_{i_{0},j_{0}})$. Therefore, $\tilde{a}_{bw}^{*}=\tilde{a}_{bi_{0}}^{*}\times\tilde{a}_{i_{0}w}^{*}=(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})$. Hence the proof. ∎
Using equations (34) and (18a) in equation (20), we get optimal interval-weights.
@@
Let $\epsilon_{i}$ and $\epsilon_{i,j}$ be as in equation (13) and (16) respectively, and let $\eta^{*}$ be the optimal objective value of problem (7). Then the following statements hold.
@@
    If $\eta^{*}=\epsilon_{i_{0}}$ for some $i_{0}\in D_{1}$, then $(\tilde{A}_{b}^{*},\tilde{A}_{w}^{*})$ given in equation (25) is the only best optimally modified PCS.
@@
    If $\eta^{*}=\epsilon_{j_{0}}$ for some $j_{0}\in D_{2}$, then $(\tilde{A}_{b}^{*},\tilde{A}_{w}^{*})$ given in equation (26) is the only best optimally modified PCS.
@@
    If $\eta^{*}=\epsilon_{i_{0},j_{0}}$ for some $i_{0},j_{0}\in D$, then $(\tilde{A}_{b}^{*},\tilde{A}_{w}^{*})$ given in equation (27) is the only best optimally modified PCS.
@@
First, assume that $\eta^{*}=\epsilon_{i_{0}}$ for some $i_{0}\in D_{1}$. By Proposition 2, $\tilde{a}_{bw}^{*}=\tilde{a}_{bw}^{\prime*}=a_{bw}-\epsilon_{i_{0}}$. Therefore, $|\tilde{a}_{bw}^{*}-a_{bw}|=|\tilde{a}_{bw}^{\prime*}-a_{bw}|$.
@@
Now, assume that $\eta^{*}=\epsilon_{i_{0},j_{0}}$ for some $i_{0},j_{0}\in D$. By Proposition 2, $\tilde{a}_{bw}^{*}=\tilde{a}_{bw}^{\prime*}=(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})$. So, $|\tilde{a}_{bw}^{*}-a_{bw}|=|\tilde{a}_{bw}^{\prime*}-a_{bw}|$.
@@
Using the best optimally modified PCS, the best optimal weight set is determined via equation (1), yielding the final resultant weights.
4.1.2 Calculation of Weights in the Presence of Multiple Best or Worst Criteria
Let $C=\{c_{1},c_{2},\ldots,c_{n}\}$ be the set of decision criteria with $c_{b_{1}},c_{b_{2}},\ldots,c_{b_{n_{1}}}$ as the best and $c_{w_{1}},c_{w_{2}},\ldots,$ $c_{w_{n_{2}}}$ as the worst criteria, and let $(A_{b},A_{w})$ be a PCS. Therefore,
$a_{b_{1}i}=a_{b_{2}i}=\ldots=a_{b_{n_{1}}i}=a_{bi}\text{ (say)},\quad a_{iw_{1}}=a_{iw_{2}}=\ldots=a_{iw_{n_{2}}}=a_{iw}\text{ (say)},$
$a_{b_{1}w_{1}}=\ldots a_{b_{1}w_{n_{2}}}=a_{b_{2}w_{1}}=\ldots a_{b_{2}w_{n_{2}}}=\ldots=a_{b_{n_{1}}w_{1}}=\ldots a_{b_{n_{1}}w_{n_{2}}}=a_{bi}\text{ (say)}$
for $i\in D^{\prime}$, where $D^{\prime}=C\setminus\{c_{b_{1}},\ldots,c_{b_{n_{1}}},c_{w_{1}}\ldots,c_{w_{n_{2}}}\}$. Whenever there is no ambiguity, we simply use the notation $D^{\prime}=\{1,\ldots,n\}\setminus\{b_{1},\ldots,b_{n_{1}},w_{1},\ldots,w_{n_{2}}\}$.
To ensure no weight difference for criteria having equal preference, instead of considering the system of equations (1), we consider the system of equations
@@ 4.1.2 Calculation of Weights in the Presence of Multiple Best or Worst Criteria
$\begin{split}&\frac{w_{b_{1}}}{w_{i}}=a_{bi},\quad\frac{w_{i}}{w_{w_{1}}}=a_{iw},\quad\frac{w_{b_{1}}}{w_{w_{1}}}=a_{bw},\quad i\in D^{\prime},\\ &w_{b_{1}}=w_{b_{2}}=\ldots=w_{b_{n_{1}}},\quad w_{w_{1}}=w_{w_{2}}=\ldots=w_{w_{n_{2}}},\quad w_{1}+w_{2}+\ldots+w_{n}=1.\end{split}$ (36)
@@
Each optimal solution of problem (37), along with $w_{b_{1}}^{*}=w_{b_{2}}^{*}=\ldots=w_{b_{n_{1}}}^{*}$ and $w_{w_{1}}^{*}=w_{w_{2}}^{*}=\ldots=w_{w_{n_{2}}}^{*}$, gives $W^{*}=\{w_{1}^{*},w_{2}^{*},\ldots,w_{n}^{*}\}$ is an optimal weight set and $\epsilon^{*}$ is a measurement of its accuracy. Since $\epsilon^{*}$ is also the optimal objective value, it remains the same for all $W^{*}$.
@@
Let $\epsilon_{i}$ and $\epsilon_{i,j}$ be as in equation (13) and (16) respectively, and let $\epsilon^{*}$ and $\eta^{*}$ be the optimal objective value of problems (37) and (40). Then
@@
The result can be proven by replicating the arguments of Theorem 2 and Proposition 2 and is thus omitted. ∎
@@
Let $\epsilon_{i}$ and $\epsilon_{i,j}$ be as in equation (13) and (16) respectively, and let $\epsilon^{*}$ and $\eta^{*}$ be the optimal objective value of problems (37) and (40). Then the following statements hold.
@@
    If $\eta^{*}=\epsilon_{i_{0}}$ for some $i_{0}\in D_{1}$, then $(\tilde{A}_{b},\tilde{A}_{w})$ defined as
@@
    If $\eta^{*}=\epsilon_{j_{0}}$ for some $j_{0}\in D_{2}$, then $(\tilde{A}_{b},\tilde{A}_{w})$ defined as
@@
    If $\eta^{*}=\epsilon_{i_{0},j_{0}}$ for some $i_{0},j_{0}\in D$, then $(\tilde{A}_{b},\tilde{A}_{w})$ defined as
@@
Using the best optimally modified PCS, the best optimal weight set is determined via equation (36), yielding the final resultant weights.
@@ 4.2 Consistency Analysis
In this subsection, our aim is to derive analytical expressions for CI and CR. From equation (33), it follows that CI can be obtained by finding the maximum possible values of $\epsilon_{i}$ and $\epsilon_{i,j}$.
@@
From equation (10), we have $(a_{bi}+\epsilon_{i})\times(a_{iw}+\epsilon_{i})=a_{bw}-\epsilon_{i}$ and $(a_{bj}+\epsilon_{j})\times(a_{jw}+\epsilon_{j})=a_{bw}-\epsilon_{j}$. Using these in equation (47), we get $a_{bw}-\epsilon_{i}<a_{bw}-\epsilon_{j}$, which is contradiction as $\epsilon_{i}<\epsilon_{j}$ implies $a_{bw}-\epsilon_{j}<a_{bw}-\epsilon_{i}$. Thus, $\epsilon_{j}\leq\epsilon_{i}$. Hence the proof. ∎
@@
The proofs of Proposition 4 and Proposition 5 are similar to the proof of Proposition 3 and thus omitted.
Proposition 3 implies that for any scale satisfying $a_{bi},a_{iw}\geq 1$ for all $i\in C$, the maximum possible value of $\epsilon_{i}$, where $i\in D_{1}$, is attained when $a_{bi}=a_{iw}=1$ for some $i\in D_{1}$. So,
$\max\{\epsilon_{i}:i\in D_{1}\}=\frac{-3+\sqrt{4a_{bw}+5}}{2}.$
Proposition 4 shows that for any scale satisfying $a_{bw}\geq 1$, the maximum possible value of $\epsilon_{i}$, where $i\in D_{2}$, is attained when $a_{bi}=a_{iw}=a_{bw}$ for some $i\in D_{2}$. So,
$\max\{\epsilon_{i}:i\in D_{2}\}=\frac{2a_{bw}+1-\sqrt{8a_{bw}+1}}{2}.$
Proposition 5 suggests that for any scale satisfying $a_{bi},a_{iw}\geq 1$ for all $i\in C$, the maximum possible value of $\epsilon_{i,j}$ is attained when $a_{bi}=a_{iw}=1$ for some $i\in D_{1}$ and $a_{bj}=a_{jw}=a_{bw}$ for some $j\in D_{2}$. So,
$\max\{\epsilon_{i,j}:i,j\in D\}=\frac{a_{bw}^{2}-1}{2a_{bw}+2}.$
From the above discussion, it follows that
$\text{CI}_{a_{bw}}=\max\biggl{\{}\frac{-3+\sqrt{4a_{bw}+5}}{2},\frac{2a_{bw}+1-\sqrt{8a_{bw}+1}}{2},\frac{a_{bw}^{2}-1}{2a_{bw}+2}\biggr{\}}.$
@@
$f(x)=\frac{2a_{bw}+1-\sqrt{8a_{bw}+1}}{2}-\frac{-3+\sqrt{4a_{bw}+5}}{2},\ x\in[1,\infty).$
@@
$\frac{2a_{bw}+1-\sqrt{8a_{bw}+1}}{2}\geq\frac{-3+\sqrt{4a_{bw}+5}}{2},$
@@
$\text{CI}_{a_{bw}}=\max\biggl{\{}\frac{2a_{bw}+1-\sqrt{8a_{bw}+1}}{2},\frac{a_{bw}^{2}-1}{2a_{bw}+2}\biggr{\}},$ (48)
which is analytical expression for CI. The values of CI for some established scales are given in Table 9.
Table 9: The values of CI for some scales
@@
22 $0.5$ $1.2222$ $0.1111$ $\sqrt{2}$ $0.2071$ $1.1257$ $0.0629$
33 11 $1.5$ $0.25$ 22 $0.5$ $1.2715$ $0.1358$
44 $1.6277$ $1.8571$ $0.4286$ $2\sqrt{2}$ $0.9142$ $1.4470$ $0.2235$
55 $2.2984$ $2.3333$ $0.6667$ 44 $1.6277$ $1.6684$ $0.3342$
66 33 33 11 $4\sqrt{2}$ $2.7563$ $1.9670$ $0.4835$
77 $3.7250$ 44 $1.6277$ 88 $4.4688$ $2.4142$ $0.7071$
88 $4.4688$ $5.6667$ $2.7633$ $8\sqrt{2}$ $7.0307$ $3.2289$ $1.1390$
99 $5.2279$ 99 $5.2279$ $16$ $10.8211$ $5.8284$ $2.8778$
Now, substituting the expressions for $\epsilon^{*}$ and CI from equations (33) and (48), respectively, into equation (6), we get
$\text{CR}=\frac{\max\left\{\bigg{|}\frac{a_{bi}+a_{iw}+1-\sqrt{(a_{bi}+a_{iw}+1)^{2}-4(a_{bi}\times a_{iw}-a_{bw})}}{2}\bigg{|},\bigg{|}\frac{a_{bi}\times a_{iw}-a_{bj}\times a_{jw}}{a_{bi}+a_{iw}+a_{bj}+a_{jw}}\bigg{|}:i,j\in D\right\}}{\max\biggl{\{}\frac{2a_{bw}+1-\sqrt{8a_{bw}+1}}{2},\frac{a_{bw}^{2}-1}{2a_{bw}+2}\biggr{\}}},$ (49)
which is analytical expression for CR.
Some important points related to the analytical expressions for CI and CR are as follows.
@@
    A consistency indicator that provides instant feedback to DMs regarding judgment inconsistencies improves the effectiveness of an MCDM method [26]. Equation (6) represents an output-based formulation of CR that lacks this ability and can only provide feedback after completing the entire calculation process. This limitation is addressed by the analytical expression of CR given in equation (49), which represents an input-based formulation of CR.
@@
    For a consistency indicator to demonstrate reasonable behavior, it must satisfy some specific properties [7, 23]. Proposition 6 outlines several of these properties without proof, as their proofs are analogous to those in [26, Proposition 1].

### v2

@@ 4.1.1 Unique Best and Worst Criterion
Now, assume that $\eta^{*}=\epsilon_{i_{0},j_{0}}\text{ for some }i_{0},j_{0}\in D$. As discussed in Proposition 1, we get $0\leq\eta_{bi_{0}},\eta_{i_{0}w},\eta_{bj_{0}},\eta_{j_{0}w}\leq\eta^{*}=\epsilon_{i_{0},j_{0}}$ such that $\tilde{a}_{bi_{0}}^{*}\in\{a_{bi_{0}}-\eta_{bi_{0}},a_{bi_{0}}+\eta_{bi_{0}}\}$, $\tilde{a}_{i_{0}w}^{*}\in\{a_{i_{0}w}-\eta_{i_{0}w},a_{i_{0}w}+\eta_{i_{0}w}\}$, $\tilde{a}_{bj_{0}}^{*}\in\{a_{bj_{0}}-\eta_{bj_{0}},a_{bj_{0}}+\eta_{bj_{0}}\}$, and $\tilde{a}_{j_{0}w}^{*}\in\{a_{j_{0}w}-\eta_{j_{0}w},a_{j_{0}w}+\eta_{j_{0}w}\}$. Since $(\tilde{A}_{b}^{*},\tilde{A}_{w}^{*})$ is consistent, we have $\tilde{a}_{bi_{0}}^{*}\times\tilde{a}_{i_{0}w}^{*}=\tilde{a}_{bj_{0}}^{*}\times\tilde{a}_{j_{0}w}^{*}$. Now, it is easy to observe that if any one of $\eta_{bi_{0}}$, $\eta_{i_{0}w}$, $\eta_{bj_{0}}$ or $\eta_{j_{0}w}$ is strictly less than $\epsilon_{i_{0},j_{0}}$, then at least one of the remaining three must be strictly greater than $\epsilon_{i_{0},j_{0}}$, which is not possible. This gives $\eta_{bi_{0}}=\eta_{i_{0}w}=\eta_{bj_{0}}=\eta_{j_{0}w}=\epsilon_{i_{0},j_{0}}$. It is clear that the only combination among the sixteen possible values of $(\tilde{a}_{bi_{0}}^{*},\tilde{a}_{i_{0}w}^{*},\tilde{a}_{bj_{0}}^{*},\tilde{a}_{j_{0}w}^{*})$ that satisfies $\tilde{a}_{bi_{0}}^{*}\times\tilde{a}_{i_{0}w}^{*}=\tilde{a}_{bj_{0}}^{*}\times\tilde{a}_{j_{0}w}^{*}$ is $(a_{bi_{0}}+\epsilon_{i_{0},j_{0}},a_{i_{0}w}+\epsilon_{i_{0},j_{0}},a_{bj_{0}}-\epsilon_{i_{0},j_{0}},a_{j_{0}w}-\epsilon_{i_{0},j_{0}})$. Therefore, $\tilde{a}_{bw}^{*}=\tilde{a}_{bi_{0}}^{*}\times\tilde{a}_{i_{0}w}^{*}=(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})$. Hence the proof. ∎
Proposition 2 implies that for any $(\tilde{A}_{b}^{*},\tilde{A}_{w}^{*})$, the value of $\tilde{a}_{bw}^{*}$ remains unchanged.
Using equations (33) and (34) in equation (20), we get optimal interval-weights.
@@
Let $\epsilon_{i}$ and $\epsilon_{i,j}$ be as in equations (13) and (16) respectively, and let $\epsilon^{*}$ be the optimal objective value of problem (3). Then the following statements hold.
@@
    If $\epsilon^{*}=\epsilon_{i_{0}}$ for some $i_{0}\in D_{1}$, then $(\tilde{A}_{b}^{*},\tilde{A}_{w}^{*})$ given in equation (25) is the only best optimally modified PCS.
@@
    If $\epsilon^{*}=\epsilon_{j_{0}}$ for some $j_{0}\in D_{2}$, then $(\tilde{A}_{b}^{*},\tilde{A}_{w}^{*})$ given in equation (26) is the only best optimally modified PCS.
@@
    If $\epsilon^{*}=\epsilon_{i_{0},j_{0}}$ for some $i_{0},j_{0}\in D$, then $(\tilde{A}_{b}^{*},\tilde{A}_{w}^{*})$ given in equation (27) is the only best optimally modified PCS.
@@
First, assume that $\epsilon^{*}(=\eta^{*})=\epsilon_{i_{0}}$ for some $i_{0}\in D_{1}$. By Proposition 2, $\tilde{a}_{bw}^{*}=\tilde{a}_{bw}^{\prime*}=a_{bw}-\epsilon_{i_{0}}$. Therefore, $|\tilde{a}_{bw}^{*}-a_{bw}|=|\tilde{a}_{bw}^{\prime*}-a_{bw}|$.
@@
Now, assume that $\epsilon^{*}(=\eta^{*})=\epsilon_{i_{0},j_{0}}$ for some $i_{0},j_{0}\in D$. By Proposition 2, $\tilde{a}_{bw}^{*}=\tilde{a}_{bw}^{\prime*}=(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})$. So, $|\tilde{a}_{bw}^{*}-a_{bw}|=|\tilde{a}_{bw}^{\prime*}-a_{bw}|$.
@@
Using this best optimally modified PCS, the best optimal weight set is obtained using equation (8).
4.1.2 Multiple Best or Worst Criteria
In cases with multiple best/worst criteria, rather than arbitrarily designating one as the best or worst, we modify the original optimization model to assign equal weights to all best criteria and equal weights to all worst criteria.
Let $c_{b_{1}},c_{b_{2}},\ldots,c_{b_{n_{1}}}$ be the best and $c_{w_{1}},c_{w_{2}},\ldots,$ $c_{w_{n_{2}}}$ be the worst criteria. Therefore,
$a_{b_{1}i}=a_{b_{2}i}=\ldots=a_{b_{n_{1}}i}=a_{bi}\text{ (say)},\quad\quad a_{iw_{1}}=a_{iw_{2}}=\ldots=a_{iw_{n_{2}}}=a_{iw}\text{ (say)},$
$a_{b_{1}w_{1}}=\ldots=a_{b_{1}w_{n_{2}}}=a_{b_{2}w_{1}}=\ldots=a_{b_{2}w_{n_{2}}}=\ldots=a_{b_{n_{1}}w_{1}}=\ldots=a_{b_{n_{1}}w_{n_{2}}}=a_{bi}\text{ (say)}$
for $i\in D^{\prime}$, where $D^{\prime}=C\setminus\{b_{1},\ldots,b_{n_{1}},w_{1},\ldots,w_{n_{2}}\}$.
To ensure $w_{b_{1}}=w_{b_{2}}=\ldots=w_{b_{n_{1}}}$ and $w_{w_{1}}=w_{w_{2}}=\ldots=w_{w_{n_{2}}}$, instead of considering the system of equations (1), consider the system of equations
@@ 4.1.2 Multiple Best or Worst Criteria
$\begin{split}&\frac{w_{b_{1}}}{w_{i}}=a_{bi},\quad\frac{w_{i}}{w_{w_{1}}}=a_{iw},\quad\frac{w_{b_{1}}}{w_{w_{1}}}=a_{bw},\quad i\in D^{\prime},\\ &w_{b_{1}}=w_{b_{2}}=\ldots=w_{b_{n_{1}}},\quad w_{w_{1}}=w_{w_{2}}=\ldots=w_{w_{n_{2}}},\\ &n_{1}w_{b_{1}}+n_{2}w_{w_{1}}+\sum_{k\in D^{\prime}}w_{k}=1.\end{split}$ (36)
@@
Each optimal solution of problem (37), along with $w_{b_{1}}^{*}=w_{b_{2}}^{*}=\ldots=w_{b_{n_{1}}}^{*}$ and $w_{w_{1}}^{*}=w_{w_{2}}^{*}=\ldots=w_{w_{n_{2}}}^{*}$, gives an optimal weight set $W^{*}=\{w_{1}^{*},w_{2}^{*},\ldots,w_{n}^{*}\}$, and $\epsilon^{*}$ is a measurement of the accuracy of this weight set. Since $\epsilon^{*}$ is also the optimal objective value, it remains the same for all $W^{*}$.
@@
Let $\epsilon_{i}$ and $\epsilon_{i,j}$ be as in equations (13) and (16) respectively, and let $\epsilon^{*}$ and $\eta^{*}$ be the optimal objective values of problems (37) and (40). Then
@@
The result can be proven by replicating the arguments of Theorem 2 and Proposition 2, and is thus omitted. ∎
@@
Equations (33) and (41) imply that for a PCS involving multiple best/worst criteria, both optimization models (3) and (37) have the same optimal objective values. This indicates that optimization model (37) selects the optimal weight sets in which all best criteria share equal weights and all worst criteria share equal weights.
@@
Let $\epsilon_{i}$ and $\epsilon_{i,j}$ be as in equations (13) and (16) respectively, and let $\epsilon^{*}$ be the optimal objective value of problem (37). Then the following statements hold.
@@
    If $\epsilon^{*}=\epsilon_{i_{0}}$ for some $i_{0}\in D_{1}$, then $(\tilde{A}_{b},\tilde{A}_{w})$ defined as
@@
    If $\epsilon^{*}=\epsilon_{j_{0}}$ for some $j_{0}\in D_{2}$, then $(\tilde{A}_{b},\tilde{A}_{w})$ defined as
@@
    If $\epsilon^{*}=\epsilon_{i_{0},j_{0}}$ for some $i_{0},j_{0}\in D$, then $(\tilde{A}_{b},\tilde{A}_{w})$ defined as
@@
Using the best optimally modified PCS, the best optimal weight set is determined using equation (8).
@@ 4.2 Consistency Analysis
This subsection derives analytical expressions for the CI and CR.
@@
From equation (10), we have $(a_{bi}+\epsilon_{i})\times(a_{iw}+\epsilon_{i})=a_{bw}-\epsilon_{i}$ and $(a_{bj}+\epsilon_{j})\times(a_{jw}+\epsilon_{j})=a_{bw}-\epsilon_{j}$. Therefore, by equation (47), we get $a_{bw}-\epsilon_{i}<a_{bw}-\epsilon_{j}$, which is contradiction as $\epsilon_{i}<\epsilon_{j}$ implies $a_{bw}-\epsilon_{j}<a_{bw}-\epsilon_{i}$. Thus, $\epsilon_{j}\leq\epsilon_{i}$. Hence the proof. ∎
Proposition 3, along with the fact that $a_{bi},a_{iw}\geq 1$ for all $i\in D_{1}$, implies that for $(A_{b},A_{w})$ with $n\geq 3$,
$\max\{\epsilon_{i}:i\in D_{1}\}=\frac{-3+\sqrt{4a_{bw}+5}}{2}.$ (48)
@@
Proof.
The proof is similar to the proof of Proposition 3, and thus omitted. ∎
Proposition 4, along with the fact that $1\leq a_{bi},a_{iw}\leq a_{bw}$ for all $i\in D_{2}$, implies that for $(A_{b},A_{w})$ with $n\geq 3$,
$\max\{\epsilon_{i}:i\in D_{2}\}=\frac{2a_{bw}+1-\sqrt{8a_{bw}+1}}{2}.$ (49)
@@
Proof.
The proof is similar to the proof of Proposition 3, and thus omitted. ∎
Proposition 5, along with the fact that $1\leq a_{bi},a_{iw}\leq a_{bw}$ for all $i\in D$, implies that for $(A_{b},A_{w})$ with $n\geq 4$,
$\max\{\epsilon_{j,k}:j\in D_{1},k\in D_{2}\}=\frac{a_{bw}^{2}-1}{2a_{bw}+2}.$ (50)
Using equations (48), (49), and (50) in equation (33), we get
$\text{CI}_{a_{bw}}(n)=\begin{cases}\max\biggl\{\frac{-3+\sqrt{4a_{bw}+5}}{2},\frac{2a_{bw}+1-\sqrt{8a_{bw}+1}}{2}\biggr\}\quad\quad\quad\quad\ \text{if }n=3,\\ \max\biggl\{\frac{-3+\sqrt{4a_{bw}+5}}{2},\frac{2a_{bw}+1-\sqrt{8a_{bw}+1}}{2},\frac{a_{bw}^{2}-1}{2a_{bw}+2}\biggr\}\quad\text{if }n\geq 4.\end{cases}$
@@
$f(x)=\frac{2a_{bw}+1-\sqrt{8a_{bw}+1}}{2}-\frac{-3+\sqrt{4a_{bw}+5}}{2},\quad x\in[1,\infty).$
@@
$\frac{2a_{bw}+1-\sqrt{8a_{bw}+1}}{2}\geq\frac{-3+\sqrt{4a_{bw}+5}}{2},$ (51)
@@
$\text{CI}_{a_{bw}}(n)=\begin{cases}\frac{2a_{bw}+1-\sqrt{8a_{bw}+1}}{2}\quad\quad\quad\ \quad\quad\quad\quad\ \text{if }n=3,\\ \max\biggl\{\frac{2a_{bw}+1-\sqrt{8a_{bw}+1}}{2},\frac{a_{bw}^{2}-1}{2a_{bw}+2}\biggr\}\quad\text{if }n\geq 4.\end{cases}$ (52)
equation (52) is an analytical expression for the CI. Table 9 presents the values of CI for the scales given in Table 1.
Table 9: The values of CI for the scales listed in Table 1
@@
$n=3$ $n\geq 4$ $n=3$ $n\geq 4$ $n=3$ $n\geq 4$ $n=3$ $n\geq 4$
22 $0.4384$ $0.5$ $1.2222$ $0.0807$ $0.1111$ $\sqrt{2}$ $0.1597$ $0.2071$ $1.1257$ $0.0441$ $0.0629$
33 11 11 $1.5$ $0.1972$ $0.25$ 22 $0.4384$ $0.5$ $1.2715$ $0.1003$ $0.1358$
44 $1.6277$ $1.6277$ $1.8571$ $0.3661$ $0.4286$ $2\sqrt{2}$ $0.8980$ $0.9142$ $1.4470$ $0.1739$ $0.2235$
55 $2.2984$ $2.2984$ $2.3333$ $0.6160$ $0.6667$ 44 $1.6277$ $1.6277$ $1.6684$ $0.2745$ $0.3342$
66 33 33 33 11 11 $4\sqrt{2}$ $2.7563$ $2.7563$ $1.9670$ $0.4215$ $0.4835$
77 $3.7250$ $3.7250$ 44 $1.6277$ $1.6277$ 88 $4.4688$ $4.4688$ $2.4142$ $0.6607$ $0.7071$
88 $4.4688$ $4.4688$ $5.6667$ $2.7633$ $2.7633$ $8\sqrt{2}$ $7.0307$ $7.0307$ $3.2289$ $1.1390$ $1.1390$
99 $5.2279$ $5.2279$ 99 $5.2279$ $5.2279$ $16$ $10.8211$ $10.8211$ $5.8284$ $2.8778$ $2.8778$
By substituting the expressions for $\epsilon^{*}$ from equation (33) and CI from equation (52) into equation (6), we get the analytical expression for CR.
Some key points regarding the CI and CR are as follows.
@@
    Since optimization models (3) and (37) produce the same optimal objective values for any PCS, equation (52) remains valid even when multiple best/worst criteria are present.
@@
    For a consistency indicator to demonstrate reasonable behavior, it must satisfy some specific properties [31, 30, 7]. Proposition 6 outlines several of these properties without proof, as their proofs are analogous to those in [34, Proposition 1].
@@
Fig. 1 shows the process for calculating optimal weights and analyzing consistency when a unique best and worst criterion exist, whereas Fig. 2 shows the corresponding process in the presence of multiple best/worst criteria.
Figure 1: Workflow for weight computation and consistency analysis (unique best and worst criterion)
Figure 2: Workflow for weight computation and consistency analysis (multiple best/worst criteria)